### v1

@@ Abstract
This study examines the relationship between upward negative stepped leader pulses in upward positive lightning and preliminary breakdown pulses (PBPs) in downward negative lightning discharges. Through analysis of simultaneous channel-base current and electric field data from the Säntis tower in Switzerland, we found notable similarities between Category A and B upward positive lightning pulses and the “Classical” and “Narrow” PBPs observed in downward negative flashes. Statistical comparisons reveal correlations between electric field and current parameters for Category A pulses, supporting the field–current relationship for preliminary breakdown proposed in previous studies. These results suggest common physical mechanisms underlying both upward and downward lightning initiation processes, providing valuable insights into lightning initiation that could not be obtained from conventional field measurements alone. Furthermore, high-speed camera footage revealed that Category B pulses can be produced by a downward-propagating recoil leader. As a whole, these findings demonstrate that detailed observations of upward lightning can offer valuable insight into the complex processes underlying lightning initiation and propagation.
@@
The initiation and early development of lightning discharges remain among the most challenging and incompletely understood aspects of atmospheric electricity research (e.g., [4]). Despite significant advancements in observation techniques, the fundamental mechanisms underlying lightning initiation continue to generate considerable scientific debate.
Numerous studies have investigated the initiation of downward negative flashes using various measurement techniques, including electric field sensors, high-speed cameras (HSC), and very high frequency (VHF) interferometry (e.g., [17, 12, 10]). The seminal work by [3] proposed a three-phase mechanism—Breakdown, Intermediate, and Leader (BIL)—that precedes the first return stroke in downward negative flashes. They hypothesized that the initial breakdown (Stage B) differs fundamentally from the stepped leader process (Stage L). This hypothesis has been both supported and challenged by subsequent research, with [17] suggesting, based on VHF observations, that similar discharge processes might occur in both the breakdown and leader stages.
@@
An earlier investigation by [1] identified parallels between the initial stage of upward negative leaders (UNLs) and the preliminary breakdown phase in downward negative flashes. They observed two distinct types of electric field pulses associated with UNLs, which they labelled “Category A” and “Category B” pulses, respectively, remarkably similar to the “Classical” and “Narrow” preliminary breakdown pulses (PBPs) identified in downward negative flashes.
In this study, we leveraged the unique measurement capabilities of the Säntis Lightning Research Facility in Switzerland to analyze simultaneous data of channel-base current, vertical electric field, and (when available) HSC recordings of the initial stage of upward positive flashes (UPFs) initiated from the Säntis tower. By performing a comparative statistical analysis of these measurements with observations of PBPs recorded in downward negative leaders reported in the literature, we aim to establish whether these similarities indicate shared physical mechanisms, potentially providing new insights into the lightning initiation process.
The paper is organized as follows: Section 2 presents the measurement results, including the different pulse type classifications and their statistical characteristics; Section 3 provides a comparative analysis between leader pulses upward lightning and PBPs in downward flashes, and examines the validity of existing models; Section 4 details the methodology employed in this study; and Section 5 summarizes our key findings and discusses their implications for lightning research.
2 Measurement results
@@ 2.1 Experimental Data
Our dataset consists of observations from ten UPFs (eight Type 2 and two Type 1 according to the classification of [19]) initiated from the Säntis tower between 2014 and 2021. For each flash, we collected current measurements in the Tower, vertical electric field measurements at a distance of 14.7 km, and, in some cases, high-speed video footage.
Table 1 provides detailed information about each flash, including their identifier, timestamp, type classification, and, in addition to the specific data available, whether or not there was any prior lightning activity in a 30 km radius in the 3 seconds prior to initiation (as confirmed by the EUCLID Lightning Location System; see [22]). We included the information on prior lightning activity because previous studies have demonstrated the impact of preceding nearby lightning on upward leader formation (e.g., [24]). See Section 4 for more methodology details.
@@
Type111According to the Romero et al. 2013’s classification. |
1 |
@@
Type 1 and Type 2 UPFs, as defined by [19], are distinguished by the presence (Type 1) or absence (Type 2) of a large unipolar return stroke-like current pulse following the upward negative stepped leader phase. However, both types share the same physical mechanism and characteristics in their early stage, beginning with an UNL initiated at the tip of the strike object, which manifests as a 100 millisecond-scale waveform with large, oscillatory pulse trains (e.g., [15]).
It should be noted that flashes 7, 8 and 9 occurred during the Laser Lightning Rod project presented in [6] (therein labelled L1, L2 and L3, respectively), while the laser was on. The guiding effect discussed therein was observed over the first 50 m of propagation (see their Fig. 2), but with no clear evidence of laser-induced lightning initiation. The presence of the laser beam does not have any obvious effect on the pulse characteristics analysed in this study (i.e., no consistent differences with significant confidence observed), though we leave the door open to this possibility as the number of observed “laser-guided” flashes was too small to draw definitive conclusions. See [14] for further discussion.
@@ 2.2 Pulse Categories and Characteristics
Through analysis of the simultaneous current and electric field measurements for over 70 pulses, we observe the two distinct categories of pulses associated with upward negative stepped leaders, as defined by [1]:
@@ 2.2.1 Category A Pulses
Category A pulses are characterized by unipolar or bipolar vertical electric field signatures, the former typically negative and the latter typically with an initial positive half-cycle111Approximately 12% exhibited inverted polarity with a negative initial half-cycle, a phenomenon also noted in PBPs of downward negative leaders [13]., correlated with negative unipolar current pulses at the tower.222We define a negative current as negative charge being transferred upward. For the E-field, we use the physics sign convention.
Bipolar E-field pulses have an average duration on the order of s and 10–90% risetime on the order of s, consistent with the characteristics of “Classical” PBPs observed in downward lightning. Unipolar E-field pulses are noticeably shorter, with an average risetime and half-width of around 1.5 and 2.5 s, respectively, though it is worth noting that these pulses were observed in only one flash (#7 in Table 1), and have therefore been excluded from the primary analysis presented here until an expanded dataset can yield more reliable results.
Both unipolar and bipolar Category A electric field pulses are well correlated with negative unipolar initial continuous current (ICC) pulses associated with the stepping of the UNL. Figure 1(a) presents the overall current and electric field waveforms associated with the upward positive flash #8 listed in Table 1. The current waveform begins with an ICC, followed by a large pulse approximately 13 milliseconds after the onset of the ICC. Figure 1(b) shows a bipolar Category A pulse train occurring in the early stage of the ICC. Figure 1(c) displays the current and E-field waveforms corresponding to a single bipolar Category A pulse.
@@ 2.2.2 Category B Pulses
Category B pulses are characterized by unipolar (positive or negative) or bipolar E-field signatures with much shorter temporal characteristics than Category A pulses, exhibiting average half-widths and risetimes of less than s. More importantly, these pulses lack correlation with any significant current pulses measured at the tower, and tend to appear later in leader development. Figure 1(d) illustrates two typical Category B pulses also belonging to flash #8, showing their faster temporal characteristics compared to Category A pulses. Note that the lack of current correlation for Category B pulses suggests a different physical mechanism than Category A; see Section 3.3.2 for further discussion.
@@ 2.3 Statistical Analysis
We conducted detailed measurements of pulse characteristics as defined in Figure 2, analyzing parameters such as the 10-90% risetime, the full width at half maximum (FWHM)333Note that [1] used the term half-peak beam width (HPBW) to refer to the same parameter., and for bipolar E-field pulses, the amplitudes and durations of both half-cycles.
Pulse Type | Category A (65) | Category B (15) |
@@
Table 2 summarizes the statistical properties of both pulse categories. and refer to the arithmetic and geometric means and standard deviations, assuming Gaussian and log-normal distributions, respectively. In addition to the average risetimes of s (AM) and s (GM), bipolar Category A electric field pulses were found to have half-widths of s (AM) and s (GM) and peak amplitudes (A1E, normalised to 100 km) of 1.0 V/m (AM) and 0.8 V/m (GM). The corresponding current pulses had risetimes of s (AM) and s (GM), half-widths of s (AM) and s (GM), and peak currents of 2.2 kA (AM) and 1.5 kA (GM). The mean step lengths are based on HSC observations of the nine Category A pulses and twelve Category B pulses measured in Flash #8. These were estimated by measuring the 2D pixel length of the channel in question, and converting to meters using the tower as a reference (124 m / 37 pix). Note that Category B step lengths are more than twice those of Category A pulses, and in both cases represent a lower bound on the true 3D length. This points to a difference in physical mechanism: Category A pulses are associated with the stepping of the UNL, i.e., virgin breakdown, while Category B pulses were observed to be due to a recoil leader retracing a pre-existing channel: see Section 3.3.2 for further discussion.
Notably, we observed time-dependent characteristics in the bipolar pulse trains, with increased peak currents (, , ) and decreased maximum E-field derivatives (, , ) as the flash progressed,444, and represent the Pearson, Spearman and Kendall correlation coefficients, respectively. as shown in Figure 1(b) and Table 3 ( represents the time from initiation of the stepped leader, i.e., the onset of the ICC). In the last two columns we can see how the step length (and its vertical component ) also grow larger. The peak amplitude AE of Category B electric field pulses also appears to increase as a function of time, though a larger dataset is needed to confirm this last, moderately non-linear, relationship (, , ). We also found significant correlations between various Category A pulse parameters, including linear relationships between the temporal widths of the initial and second half-cycles (, , ) and between their peak amplitudes (, , ).
Category B pulses demonstrated a very strong linear relationship between their amplitude and maximum derivative (, , ), yielding a minimum characteristic risetime of AE/ = 0.3 s (with a coefficient of determination ), about 1/2 the measured R1E.
@@
0 | 3.480.02 | 5.480.02 | 1.920.05 | 9.30.3 | 2.40.1 | 0.90.1 | 0.400.05 | 5.90.5 | 144 | 103 |
46 | 3.38 | 5.88 | 3.00 | 8.0 | 1.1 | 0.8 | 0.55 | 9.2 | 17 | 10 |
145 | 11.78 | 8.38 | 0.64 | 2.9 | 0.5 | 0.9 | 0.14 | 5.1 | 22 | 20 |
267 | 3.98 | 5.88 | 0.58 | 1.2 | 2.9 | 2.4 | 0.11 | 3.4 | 22 | 20 |
455 | 3.88 | 6.38 | 0.90 | 1.1 | 1.4 | 1.8 | 0.23 | 3.8 | 24 | 20 |
631 | 4.38 | 7.18 | 0.48 | 1.1 | 1.0 | 1.8 | 0.17 | 4.9 | 36 | 27 |
817 | 4.48 | 14.18 | 0.44 | 0.7 | 2.0 | 1.3 | 0.09 | 3.7 | 34 | 20 |
989 | 10.98 | 8.08 | 0.60 | 0.9 | 2.1 | 7.8 | 0.15 | 3.2 | 36 | 31 | 34 | 20 111This “step” exhibited simultaneous extension of both branches of the plasma channel. |
5.673.06 | 7.682.65 | 1.070.86 | 3.13.2 | 1.70.8 | 2.22.2 | 0.230.15 | 6.01.5 | 268 | 207 |
3 Similarity with PBPs in downward flashes
3.1 Overview of PBPs in Downward Lightning
@@
The physical mechanisms underlying these pulses remain contested. Using high-speed video footage, [23] proposed an “initial leader” concept for Classical PBPs (in stage B) that differs from normal stepped leaders (in stage L), involving a dim linear feature moving downward, followed by impulsive breakdown at the lower end and an upward-moving brightness. They hypothesized that the initial leader pulses stop occurring because the previous initial leaders transferred enough charge to reduce the ambient electric field near the initiation point. In contrast, [2] observed that channel extension in stage B appears similar to ordinary leader extension in stage L, suggesting comparable mechanisms.
[16] observed a bimodal distribution of the stepping process, involving both long (200+ m length) and short (10+ m length) steps. The same kind of pulses observed in stage L (“Narrow” PBPs) were observed in stage B, superimposed on slower, “Classical” PB pulses. They suggested the presence of distant space leaders with their negative ends stepping downward and generating narrow pulses (like those in stage L), while their positive ends generate classical pulses when they attach to the previous main leader channel (major pulses in stage B). With the ambient electric field reduced by the descent of the leader, less distant space leaders occur, which leads to the generation of only narrow pulses shortly before the return stroke, as the leader approaches the ground.
3.2 Comparative Analysis of Pulse Characteristics

### v2

@@ Abstract
This study examines the relationship between upward negative stepped leader pulses in upward positive lightning and preliminary breakdown pulses (PBPs) in downward negative lightning discharges. Through analysis of simultaneous channel-base current and electric field data from the Säntis tower in Switzerland, we found notable similarities between the “Category A” and “B” pulses associated with the initial continuous current of upward negative leaders, and the “Classical” and “Narrow” PBPs observed in downward negative flashes. Statistical comparisons reveal correlations between measured electric field and tower current parameters for Category A pulses, supporting the field–current relationship for preliminary breakdown proposed in previous studies. These results suggest that similar physical processes may be involved in the early stages of negative leader development in both upward and downward lightning, providing valuable insights into lightning initiation that could not be obtained from conventional field measurements alone. Furthermore, high-speed camera footage revealed that Category B pulses can be produced by a downward-propagating recoil leader. As a whole, these findings demonstrate that detailed observations of upward lightning can offer valuable insight into the complex processes underlying lightning initiation and propagation.
@@
The initiation and early development of lightning discharges remain among the most challenging and incompletely understood aspects of atmospheric electricity research (e.g., [dwyer_physics_2014]). Despite significant advancements in observation techniques, the fundamental mechanisms underlying lightning initiation continue to generate considerable scientific debate.
Numerous studies have investigated the initiation of downward negative flashes using various measurement techniques, including electric field sensors, high-speed cameras (HSC), and very high frequency (VHF) interferometry (e.g., [proctor_vhf_1988, nag_analysis_2009, marshall_percentage_2014]), in both tropical and temperate thunderstorms [gomes_comparison_1998]. The seminal work by [clarence_preliminary_1957] proposed a three-phase mechanism—Breakdown, Intermediate, and Leader (BIL)—that precedes the first return stroke in downward negative flashes. They hypothesised that the initial breakdown (Stage B) differs fundamentally from the stepped leader process (Stage L). This hypothesis has been both supported and challenged by subsequent research, with [proctor_vhf_1988] suggesting, based on VHF observations, that similar discharge processes might occur in both the breakdown and leader stages.
@@
An earlier investigation by [azadifar_similarity_2018] identified parallels between the initial stage of upward negative leaders (UNLs) and the preliminary breakdown phase in downward negative flashes. They observed two distinct types of electric field pulses associated with UNLs, which they labelled “Category A” and “Category B” pulses, respectively, remarkably similar to the “Classical” and “Narrow” preliminary breakdown pulses (PBPs) identified in downward negative flashes.
In this study, we expand on their dataset and analysis, leveraging the measurement capabilities of the Säntis Lightning Research Facility in Switzerland to analyse simultaneous data of channel-base current, vertical electric field, and (when available) HSC recordings of the initial stage of upward positive flashes (UPFs) initiated from the Säntis Tower. By performing a more comprehensive comparative statistical analysis of these measurements with observations of PBPs recorded in downward negative leaders reported in the literature, we aim to establish whether these similarities are consistent with related physical processes, potentially providing new insights into the early stages of lightning leader development and possibly initiation. Moreover, the inclusion of current pulse characteristics and HSC recordings in this expanded analysis allows us to validate existing physical models for PBPs in the context of UNL pulses, further supporting the notion of a shared origin.
The paper is organised as follows: Section 2 details the methodology employed in this study; Section 3 presents the measurement results, including the different pulse type classifications and their statistical characteristics; Section 4 provides a comparative analysis between leader pulses in upward lightning and PBPs in downward flashes, and examines the validity of existing models; and Section 5 summarises our key findings and discusses their implications for lightning research.
2 Methods
The Mt. Säntis Lightning Research Facility, shown in Figure 1, is situated at 2502 m ASL in the Appenzell Alps of north-eastern Switzerland, and experiences 100 direct lightning strikes per year to its 124-meter-tall tower, which is equipped with a comprehensive current measurement system consisting of Rogowski coil and B-dot sensor pairs at two different heights: 24 and 82 meters above ground level (AGL). The system has a sampling rate of 50 MHz. Five kilometers away, atop Mt. Krönberg (1663 m ASL), is a high-speed camera (HSC) operating at 24,000 fps, with an exposure time of s and a resolution of 512512 pixels. Fifteen kilometres away, atop a building in Herisau, Switzerland (771 m ASL) with line-of-site, lies an flat-plate electric field antenna with a frequency range of 40 Hz to 40 MHz, a sampling rate of 10 MHz, and time constants of 4.2 ms (in 2014) and 8.4 ms (2021). These settings are such that the sub-millisecond pulse temporal parameters measured herein are unaffected; furthermore, we do not expect significant signal propagation effects on these time scales at this distance. This E-field signal is synchronised with the tower current signal by GPS timestamp if the antennae are functional at the time of the flash. If not, manual synchronisation can be carried out via waveform matching, with an uncertainty of s. More detailed information on the Säntis measurement system can be found in [rachidi_santis_2022].
3 Measurements
@@ 3.1 Experimental Data
Our dataset consists of observations from ten UPFs (eight Type 2 and two Type 1 according to the classification of [romero_positive_2013]) initiated from the Säntis tower between 2014 and 2021. For each flash, we collected current measurements in the Tower, vertical electric field measurements at a distance of 14.7 km, and, in one case, high-speed video footage.
Table 1 provides detailed information about each flash, including their identifier, timestamp, and, in addition to the specific data available, whether or not there was any prior lightning activity in a 30 km radius in the 3 seconds prior to initiation (as confirmed by the EUCLID Lightning Location System; see [smorgonskiy_analysis_2015]). We included the information on prior lightning activity as previous studies have demonstrated the impact of preceding nearby lightning on upward leader formation (e.g., [wang_observed_2008, sunjerga_initiation_2021]). See Section 2 for more methodology details.
@@
Prior Activity111in a 30 km radius in the 3 seconds prior to initiation. | Current (& derivative) | High-speed camera | E-field (15-km) |
| 1 |
@@
Type 1 and Type 2 UPFs, as defined by [romero_positive_2013], are distinguished by the presence (Type 1) or absence (Type 2) of a large unipolar return stroke-like current pulse following the upward negative stepped leader phase. However, both types share the same physical mechanism and characteristics in their early stage, beginning with an UNL initiated at the tip of the strike object, which manifests as a waveform lasting on the order of 100 milliseconds, which is punctuated by large, oscillatory pulse trains (e.g., [heidler_characteristics_2015, oregelchaumont_underlying_2025]). These observations are further confirmed by the statistical similarity of the current and E-field pulse characteristics: with the exception of average current amplitudes being larger by a factor of for Type 2 UPFs, all other calculated means fell within a standard deviation of each other for Type 1 and 2 UPFs. As such, both types are grouped together in the following analysis.
It should be noted that flashes 7, 8 and 9 occurred during the Laser Lightning Rod project presented in [houard_laser-guided_2023] (therein labelled L1, L2 and L3, respectively), while the laser was on. The guiding effect discussed therein was observed over the first 50 m of propagation (see their Fig. 2), but with no clear evidence of laser-induced lightning initiation. The presence of the laser beam does not have any obvious effect on the pulse characteristics analysed in this study (i.e., no consistent differences with significant confidence observed), though we leave the door open to this possibility as the number of observed “laser-guided” flashes was too small to draw definitive conclusions. See [oregel-chaumont_direct_2024] for further discussion.
@@ 3.2 Pulse Categories and Characteristics
Through analysis of the simultaneous current and electric field measurements for over 70 pulses, we observe the two distinct categories of pulses associated with upward negative stepped leaders, as defined by [azadifar_similarity_2018]:
@@ 3.2.1 Category A Pulses
Category A pulses are characterised by unipolar or bipolar vertical electric field signatures, the former typically negative and the latter typically with an initial positive half-cycle111Approximately 12% exhibited inverted polarity with a negative initial half-cycle, a phenomenon also noted in the PBPs of downward negative leaders [ogawa_initiation_1993]., correlated with negative unipolar current pulses at the tower.222We define a negative current as negative charge being transferred upward. For the E-field, we use the physics sign convention.
Bipolar E-field pulses have an average duration on the order of s and 10–90% risetime on the order of s, consistent with the characteristics of “Classical” PBPs observed in downward lightning. Unipolar E-field pulses are noticeably shorter, with an average risetime and half-width of around 1.5 and 2.5 s, respectively, comparable to values measured by [wu_upward_2020]. It is worth noting, however, that we observed these pulses in only one flash (#7 in Table 1), and have therefore been excluded from the primary analysis presented here until an expanded dataset can yield more reliable results.
Both unipolar and bipolar Category A electric field pulses are well correlated with negative unipolar initial continuous current (ICC) pulses associated with the stepping of the UNL. Figure 2(a) presents the overall current and electric field waveforms associated with the upward positive flash #8 listed in Table 1. The current waveform begins with an ICC, followed by the characteristic Type 1 main pulse approximately 13 milliseconds after the onset of the ICC. Figure 2(b) shows a bipolar Category A pulse train occurring in the early stage of the ICC. Figure 2(c) displays the current and E-field waveforms corresponding to a single bipolar Category A pulse.
@@ 3.2.2 Category B Pulses
Category B pulses are characterised by unipolar (positive or negative) or bipolar E-field signatures with much shorter temporal characteristics than Category A pulses, exhibiting average half-widths and risetimes of less than s. More importantly, these pulses lack correlation with any significant current pulses measured at the tower, and tend to appear later in leader development. Figure 2(d) illustrates two typical Category B pulses also belonging to flash #8, showing their faster temporal characteristics compared to Category A pulses. Note that the lack of current correlation for Category B pulses suggests a different physical mechanism than Category A; see Section 4.3.2 for further discussion.
4 Results
@@ 4.1 Statistical Analysis
We conducted detailed measurements of pulse characteristics as defined in Figure 3, analysing parameters such as the 10-90% risetime, the full width at half maximum (FWHM)333Note that [azadifar_similarity_2018] used the term half-peak beam width (HPBW) to refer to the same parameter., and for bipolar E-field pulses, the amplitudes and durations of both half-cycles.
| Pulse Type () | Category A (65) | Category B (15) |
@@
Table 2 summarises the statistical properties of both pulse categories, based on 65 Category A pulses and 15 Category B pulses. and refer to the arithmetic and geometric means and standard deviations, assuming Gaussian and log-normal distributions, respectively. In addition to the average risetimes of s (AM) and s (GM), bipolar Category A electric field pulses were found to have half-widths of s (AM) and s (GM) and peak amplitudes (A1E, normalised to 100 km) of 1.0 V/m (AM) and 0.8 V/m (GM). The corresponding current pulses had risetimes of s (AM) and s (GM), half-widths of s (AM) and s (GM), and peak currents of 2.2 kA (AM) and 1.5 kA (GM). The mean step lengths are based on HSC observations of the nine Category A pulses and twelve Category B pulses measured in Flash #8. These were estimated by measuring the 2D pixel length of the channel in question, and converting to meters using the tower as a reference (124 m / 37 pix). Note that Category B step lengths are more than twice those of Category A pulses, and in both cases represent a lower bound on the true 3D length. This points to a difference in physical mechanism: Category A pulses are associated with the stepping of the UNL, i.e., virgin breakdown, while Category B pulses were observed to be due to a recoil leader retracing a pre-existing channel: see Section 4.3.2 for further discussion.
Notably, we observed time-dependent characteristics in the bipolar pulse trains, with increased peak currents (, , ) and decreased maximum E-field derivatives (, , ) as the flash progressed,444, and represent the Pearson, Spearman and Kendall correlation coefficients, respectively, calculated from the aforementioned sample size of . as shown in Figure 2(b) and Table 3 ( represents the time from initiation of the stepped leader, i.e., the onset of the ICC). In the last two columns we can see how the step length (and its vertical component ) also grow larger. We also found significant correlations between various Category A pulse parameters, including linear relationships between the temporal widths of the initial and second half-cycles (, , ) and between their peak amplitudes (, , );see Figure 4.
Category B pulses demonstrated a very strong linear relationship between their amplitude and maximum derivative (, , ), yielding a minimum characteristic risetime of AE/ = 0.3 s (with a coefficient of determination ), about 1/2 the measured risetime R1E, though a larger dataset is needed to confirm this last relationship, due to the relatively small sample size of .
@@
| 0 | 3.480.02 | 5.480.02 | 1.920.05 | 9.30.3 | 2.40.1 | 0.90.1 | 0.400.05 | 5.90.5 | 144 | 103 |
| 46 | 3.38 | 5.88 | 3.00 | 8.0 | 1.1 | 0.8 | 0.55 | 9.2 | 17 | 10 |
| 145 | 11.78 | 8.38 | 0.64 | 2.9 | 0.5 | 0.9 | 0.14 | 5.1 | 22 | 20 |
| 267 | 3.98 | 5.88 | 0.58 | 1.2 | 2.9 | 2.4 | 0.11 | 3.4 | 22 | 20 |
| 455 | 3.88 | 6.38 | 0.90 | 1.1 | 1.4 | 1.8 | 0.23 | 3.8 | 24 | 20 |
| 631 | 4.38 | 7.18 | 0.48 | 1.1 | 1.0 | 1.8 | 0.17 | 4.9 | 36 | 27 |
| 817 | 4.48 | 14.18 | 0.44 | 0.7 | 2.0 | 1.3 | 0.09 | 3.7 | 34 | 20 |
| 989 | 10.98 | 8.08 | 0.60 | 0.9 | 2.1 | 7.8 | 0.15 | 3.2 | 36 | 31 | 34 | 20 111This “step” exhibited simultaneous extension of both branches of the plasma channel. |
| 5.673.06 | 7.682.65 | 1.070.86 | 3.13.2 | 1.70.8 | 2.22.2 | 0.230.15 | 6.01.5 | 268 | 207 |
4.2 Similarity with PBPs in downward flashes
4.2.1 Overview of PBPs in Downward Lightning
@@
The physical mechanisms underlying these pulses remain contested. Using high-speed video footage, [stolzenburg_luminosity_2013] proposed an “initial leader” concept for Classical PBPs (in stage B) that differs from normal stepped leaders (in stage L), involving a dim linear feature moving downward, followed by impulsive breakdown at the lower end and an upward-moving brightness. They hypothesised that the initial leader pulses stop occurring because the previous initial leaders transferred enough charge to reduce the ambient electric field near the initiation point. In contrast, [campos_visible_2013] observed that channel extension in stage B appears similar to ordinary leader extension in stage L, suggesting comparable mechanisms.
[petersen_high-speed_2015] observed a bimodal distribution of the stepping process, involving both long (200+ m length) and short (10+ m length) steps. The same kind of pulses observed in stage L (“Narrow” PBPs) were observed in stage B, superimposed on slower, “Classical” PB pulses. They suggested the presence of distant space leaders with their negative ends stepping downward and generating narrow pulses (like those in stage L), while their positive ends generate classical pulses when they attach to the previous main leader channel (major pulses in stage B). With the ambient electric field reduced by the descent of the leader, less distant space leaders occur, which leads to the generation of only narrow pulses shortly before the return stroke, as the leader approaches the ground.
4.2.2 Comparative Analysis of Pulse Characteristics